OpenStudy (anonymous):

((9-x^2)^2)^(1/3) use the chain rule to find the derivative. Can it also be written like this: (9−x^2)2/3?

OpenStudy (anonymous):

hold on almost got it

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

(4x^3-36x) (x^4-18x^2+81)^(-2/3) -------------------------------- 3

OpenStudy (anonymous):

how did you do that?

OpenStudy (anonymous):

haha it's a bit involved but i got ya

OpenStudy (anonymous):

just the first step is fine :)

OpenStudy (anonymous):

or the first two or 3 :)

OpenStudy (anonymous):

what i first did was did the FOIL method for (9-x^2)^2 so that we didn't have power to a power to a power thing and that came out to be x^4-18x^2+81

OpenStudy (anonymous):

so then you got this (x^4-18x^2+81)^(1/3)

OpenStudy (anonymous):

did you use the chain rule?

OpenStudy (anonymous):

not yet

OpenStudy (anonymous):

then you chunk the whole thing as a variable and deferenciate like normal. so we bring down the (1/3) and put that as a coeffiecent in the front and then re write the whoel thing and then subtract one from the power, just like a normal thing

OpenStudy (anonymous):

but since you did multiple variables at the same time, you have to use the chain rule

OpenStudy (anonymous):

and that involves you taking the derivative of the chunk that we just deferenciated so that was (1/3) (x^4-18x^2+81)^(-2/3) *applying chain rule* (4x^3-36x) so then you get this (1/3) (4x^3-36x) (x^4-18x^2+81)^(-2/3)

OpenStudy (anonymous):

and you're done

OpenStudy (anonymous):

thank you, something that has always confused me is negative exponents, for example, (9x+8)^(-2/3) how would you write that?

OpenStudy (anonymous):

ahh yeah it's not bad. A negative exponet means 1/(that thing) so for example x^(-2) = 1/(x^2)

OpenStudy (anonymous):

but when it' a fraction like that you alwasy have to remember (power/root) so if we had x(3/2) / 1/(SQRT(x^3))

OpenStudy (anonymous):

opps i meant x(3/2) = 1/(SQRT(x^3))

OpenStudy (anonymous):

Watch a few videos from this guy. He's really good and does a ton of examples http://www.youtube.com/watch?v=6kScLENCXLg

OpenStudy (anonymous):

thank you for all your help!

OpenStudy (anonymous):

no problem